OpenStudy (anonymous):

OK, can't figure out now how (x^(1+sqrt(2)))/(1+sqrt(2)) simplifies to (sqrt(2)-1)x^(sqrt(2)+1) Help me out here please.

OpenStudy (pokemon23):

Hey Pansi can you help me?

OpenStudy (pokemon23):

hey mert can you help me?

OpenStudy (mertsj):

Pasi, if what you have typed is correct, the exponent, 1+sqrt2 divided by 1+sqrt2 = 1 so your expression would be x^1. You must have typed something wrong.

OpenStudy (anonymous):

Clearer question, how: \[(x^{\sqrt(2)+1})/(\sqrt(2)+1)\] simplifies to: \[(\sqrt(2)-1)x^{\sqrt(2)+1}\] ?

OpenStudy (phi):

multiply top and bottom by 1-sqrt(2)

OpenStudy (phi):

\[(1+\sqrt2)(1-\sqrt2)= 1-2= -1 \] in the denominator.

OpenStudy (phi):

using (a-b)(a+b)= a^2 - b^2

OpenStudy (phi):

get rid of the -1 in the denominator by multiplying top and bottom by -1, this changes the 1-sqrt(2) to sqrt(2)-1

OpenStudy (anonymous):

Of course! Thank you very much!

OpenStudy (phi):

fyi, you always use this trick to get rid of radicals in the denominator.

OpenStudy (anonymous):

Yes, I knew it ones, but had forgotten it :) Thank you for the reminder :D

OpenStudy (anonymous):

Find the critical point(s) of \[f(x,y)=e ^{3x ^{2}}+e ^{y ^{2}}\] and check whether it is relative maxima, relative minima or a saddle point by using second partial derivative test.

OpenStudy (anonymous):

?